Two cars each complete a journey of 120 km.
The first car is driven at an average speed of x km/h.
The second car is driven at an average speed 3 km/h faster than the first car.
The first car takes 6 minutes longer to complete the journey.
Write down an equation in x and show that it simplifies to x2 + 3x – 3600 = 0.

Answer by TimothyLamb(4379)

x = speed of first car
y = speed of second car
t = time of second car
---
y = x + 3
---
s = d/t
t = d/s
---
time of first car:
t + 6/60 = 120/x
t = 120/x - 6/60
---
time of second car:
t = 120/y
t = 120/(x + 3)
---
equate times:
120/x - 6/60 = 120/(x + 3)
120/x - 120/(x + 3) = 6/60
120(x + 3)/x(x + 3) - 120x/x(x + 3) = 6/60
120(x + 3) - 120x = (6/60)x(x + 3)
120x + 360 - 120x = (6/60)xx + 3*(6/60)x
360 = (6/60)xx + 3*(6/60)x
(6/60)xx + 3*(6/60)x - 360 = 0
( (6/60)xx + 3*(6/60)x - 360 = 0 )/6
(1/60)xx + 3*(1/60)x - 60 = 0
( (1/60)xx + 3*(1/60)x - 60 = 0 )*60
(1/1)xx + 3*(1/1)x - 60*60 = 0
---
answer:
xx + 3x - 3600 = 0
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---